OpenStudy (anonymous):

cos x +cos 2x =0 use exact values in radians

OpenStudy (anonymous):

start by replacing \(\cos(2x)\) by \(2\cos^2(x)-1\) and solve the quadratic equation

OpenStudy (anonymous):

cos2x =2cos^2 x -1 cos x +cos 2x =0 cosx+2cos^2 x -1=0 let u=cosx 2u^2 +u-1=0 (u+1)(2u-1)=0 u=-1 or u=1/2 cosx=-1 or cosx=1/2 x=pi,3pi,.... or x=pi/3 or -pi/3 ....